# Adaptive Exploration for Multi-Reward Multi-Policy Evaluation

### v2

@@ 6 Numerical Results
In this section, we present the numerical results of MR-NaS, and other algorithms, on various environments with different reward sets (results consider $30$ seeds).
@@ Discussion.
The results for the first two settings are shown in Figures 3 and 4 (policy evaluation was performed using the MDP estimate $M_{t}$ at each time-step). MR-NaS achieves good accuracy on all environments. On the other hand, SF-NR and GVFExplorer have mixed performance. While SF-NR is not designed to optimize a behavior policy, we note that the exploration strategy used by GVFExplorer is similar to solving a problem akin to Equation 2, but neglects the forward equations when optimizing the behavior policy (see also Section C.3 for details). As a result, GVFExplorer tends to perform worse in environments where certain rewards are hard to obtain under a uniform policy. Lastly, we refer the reader to Section C.4 for more details, and to the README.me file in the supplementary material to reproduce the results.
@@ 7 Conclusions
In this work, we studied the problem of devising an exploration strategy for online multi-reward multi-policy evaluation, accommodating reward sets that are either finite or convex, potentially encompassing all possible rewards. Leveraging tools from Best Policy Identification, we derived an instance-dependent sample complexity lower bound for the $(\epsilon,\delta)$-PAC setting. Based on this bound, we extended MR-NaS (Russo & Vannella, 2024) to this policy-evaluation setting, and showed its asymptotic efficiency. Lastly, we compared MR-NaS against other adaptive exploration across various domains, demonstrating its efficiency.
@@ Appendix A Additional Notation
$P_{s}^{\pi}$ $P_{s}^{\pi}=P(s,\pi(s))$ Vector of transition probabilities under $\pi$ in $s$.
@@
$G^{\pi}$ $G^{\pi}=(I-\gamma P^{\pi})^{-1},G^{\pi}\in\mathbb{R}^{S\times S}$ Stationary (unnormalized) state distribution.
$\rho_{r}^{\pi}(s,s^{\prime})$ $\rho_{r}^{\pi}(s,s^{\prime})=V_{r}^{\pi}(s^{\prime})-\mathbb{E}_{\hat{s}\sim P(s,\pi(s))}[V_{r}^{\pi}(\hat{s})]$ One-step value deviation in $(s,s^{\prime})$.
@@ B.1 Alternative Models and Value Deviation
We begin by proving Proposition 4.1 and an additionel necessary (minimax) condition for a model $M^{\prime}$ to be an alternative mode.
B.1.1 Proof of Proposition 4.1 and an Additional Minimax Cecessary Condition
@@
Sufficient condition. Let $s_{0},s_{1}\in{\cal S}$ and fix a reward $r\in{\cal R}$. In the following we omit the subscript $r$ for simplicity. Define a transition function $P^{{}^{\prime}}$ so that
@@
Letting $M=\sum_{i}(\gamma P^{{}^{\prime}\pi})^{i}e_{s_{0}}$, observe that $M_{s_{0}}=e_{s_{0}}^{\top}M$ represents the (un-normalized) stationary discounted probability of reaching $s_{0}$ starting from $s_{0}$
@@
Let $p_{s_{0}}^{\pi}=P(s_{0}|s_{0},\pi(s_{0}))$ Then, we seek a value of $\delta\in(0,1)$ so that
@@
Necessary condition. As before, let $V_{M}^{\pi}$ be the value of $\pi$ in $M$ with a reward $r$ (we omit the subscript $r$ for simplicity), and similarly define $V_{M^{\prime}}^{\pi}$. If $M^{\prime}$ is confusing for some reward $r\in{\cal R}$, then we have that there exists a state $s_{0}$ such that $2\epsilon<|\Delta V^{\pi}(s_{0})|$, with $\Delta V^{\pi}(s)=V_{P}^{\pi}(s)-V_{M^{\prime}}^{\pi}(s)$. Note that the following inequalities hold for any $s$
$|\Delta V^{\pi}(s)|$ $\leq\gamma\left|P(s,\pi(s))^{\top}V_{M}^{\pi}-P_{r}^{\prime}(s,\pi(s))^{\top}V_{M^{\prime}}^{\pi}\right|,$
@@
In the following proof we provide an alternative minimax necessary condition for the existance of an alternative model for $(\pi,r)$. To derive the result, we use the following lemma.
@@
The divergence between discounted future state visitation distributions, $||d^{\pi^{\prime}}-d^{\pi}||_{1}$, is bounded by an average divergence of the policies $\pi^{\prime}$ and $\pi$: $||d^{\pi^{\prime}}-d^{\pi}||_{1}\leq\frac{2\gamma}{1-\gamma}\mathbb{E}_{s\sim d^{\pi}}[\|\bar{\pi}(s)-\bar{\pi}^{\prime}(s)\|_{TV}]$
@@
Consider a reward $r$ and a target policy $\pi$. If $M^{\prime}$ is a confusing model, i.e. $||V^{\pi}_{M^{\prime}}-V^{\pi}_{M}||_{\infty}>2\epsilon$, then there $\exists s$ such that $\mathbb{E}_{a\sim\pi(\cdot|s)}[{\rm KL}_{P|P^{\prime}}(s,a)]>\frac{2(1-\gamma)^{4}}{\gamma^{2}}\epsilon^{2}.$
@@
The proof relies on constructing an alternative MDP and a policy $\bar{\pi}$ such that we can transform the problem of evaluating two different transition function into a problem of comparing two different policies.
@@
where $d_{\bar{M}_{\bar{r}}}^{\pi}(s)$ denotes the visitation distribution induced by $\pi$ in $\bar{M}_{\bar{r}}$ starting from $s$.
@@
$\frac{(1-\gamma)^{4}}{\gamma^{2}}\epsilon\leq\|\bar{\pi}(s)-\bar{\pi}^{\prime}(s)\|_{TV}^{2}\leq\frac{1}{2}{\rm KL}(\bar{\pi}(s),\bar{\pi}^{\prime}(s))\leq\frac{1}{2}\mathbb{E}_{a\sim\pi(\cdot|s)}\left[{\rm KL}_{P|P^{\prime}}(s,a)\right].$
@@ B.1.2 Additional Results on the Value Deviation $\rho$
${\rm sp}(V_{r}^{\pi})\coloneqq\max_{s^{\prime}}V_{r}^{\pi}(s^{\prime})-\min_{s^{\prime}}V_{r}^{\pi}(s),$ (11)
@@
For any reward vector $r\in[0,1]^{S}$, states $s,s^{\prime}\in{\cal S}$, we have: (I) $|\rho_{r}^{\pi}(s,s)|\leq 1$; (II) $|\rho_{r}^{\pi}(s,s^{\prime})|\leq{\rm sp}(V_{r}^{\pi})$ and (III) $|\rho_{r}^{\pi}(s,s^{\prime})-\rho_{r}^{\pi}(s^{\prime},s)|\leq|\Delta_{r}^{\pi}(s,s^{\prime})|+\Gamma_{s,s^{\prime}}^{\pi}{\rm sp}(V_{r}^{\pi}),$ where $\Delta_{r}^{\pi}(s,s^{\prime})\coloneqq r(s,\pi(s))-r(s^{\prime},\pi(s^{\prime}))$ and $\Gamma_{s,s^{\prime}}^{\pi}\coloneqq\frac{1+\gamma}{2}\max_{s,s^{\prime}}\|P(s,\pi(s))-P(s^{\prime},\pi(s^{\prime}))\|_{1}$.
@@
Then, define ${\cal Z}(s,s^{\prime})\coloneqq\{z\in{\cal S}:P(z|s,\pi(s))\geq P(z|s^{\prime},\pi(s^{\prime}))\}$ and
@@
The second statement stems from the simple fact that $-V_{r}^{\pi}(s^{\prime})\leq-\min_{s}V_{r}^{\pi}(s)$ for any $s^{\prime}$. Hence:
@@
For any reward vector $r\in[0,1]^{S}$, we have:$\rho_{r}^{\pi}(s,s^{\prime})\in\left[-\frac{1-\lambda_{r}^{\pi}\gamma}{1-\gamma},\max\left(1,\frac{1-\lambda_{r}^{\pi}}{1-\gamma}\right)\right]$ for all $s,s^{\prime}\in{\cal S}$. Furthemore, for $\lambda_{r}^{\pi}=\gamma$ we have that $|\rho_{r}^{\pi}(s,s^{\prime})|\leq 2$ for all $s,s^{\prime}$.
@@
Second part. For the second part of the lemma let $P_{s}^{\pi}=P(s,\pi(s))$ be the vector of transition probabilities in $s$ under action $\pi(s)$. We first prove that $V_{r}^{\pi}(s)\geq\lambda_{r}^{\pi}V_{r}^{\pi}(s^{\prime})\Rightarrow(P_{s}^{\pi})^{\top}V_{r}^{\pi}\geq\lambda_{r}^{\pi}(P_{s^{\prime}}^{\pi})^{\top}V_{r}^{\pi}$ The proof is simple, and follows from the following inequalities that hold for any pair $(s,\hat{s})\in{\cal S}^{2}$:
@@
    We now prove a lower bound on $\rho_{r}^{\pi}(s,s^{\prime})$. The idea is to seek a value of $\eta\geq 0$ such that $\gamma W_{s^{\prime}}-W_{s}\geq(\gamma-1)W_{s}\eta$, which implies $\gamma W_{s^{\prime}}-W_{s}\geq-\eta$ (using that $-W_{s}\geq-1/(1-\gamma)$). Then, we find $W_{s^{\prime}}\geq\frac{(\eta\gamma-\eta+1)}{\gamma}W_{s}$. Since $W_{s^{\prime}}\geq\lambda_{r}^{\pi}W_{s}$, we can set
@@
In the following proof, we prove Proposition 4.2, which characterizes the set of rewards for which $\max_{s}\|\rho_{r}^{\pi}(s)\|_{\infty}=0$. In the proof we define the following deviation matrix
@@
where $P^{\pi}$ is a $S\times S$ matrix satisfying $(P^{\pi})_{s,s^{\prime}}=P(s^{\prime}|s,\pi(s))$. We consider the condition $\rho_{r}^{\pi}=0$ being identically 00, that is
@@ Conclusion.
The set of all vectors $r$ for which $\rho_{r}^{\pi}=0$ is precisely the one-dimensional subspace $\{\alpha{\bf 1}:\alpha\in\mathbb{R}\}$. Thus, $r$ must be a scalar multiple of ${\bf 1}$. ∎
@@
Let ${\rm diag}(\rho_{r}^{\pi})=\begin{bmatrix}\rho_{r}^{\pi}(s_{1},s_{1})&\dots\rho_{R}^{\pi}(s_{S},s_{S})\end{bmatrix}^{\top}$. We have that ${\rm diag}(\rho_{r}^{\pi})=(I-P^{\pi})(I-\gamma P^{\pi})^{-1}r.$ (13)
@@
We now consider a result that permits to understand when $\max_{s}\rho_{r}^{\pi}(s,s)=0$ for any state $s$, which is the diagonal of $\rho_{r}^{\pi}$. In the result, we consider a partition of the set of recurrent states of $P^{\pi}$ into $m$ disjoint closed irreducible sets $C_{1},\dots,C_{m}$. Then, define $\tau_{i}=\inf\{t:s_{t}\in C_{i}\}$ be the hitting time of $C_{i}$, that is, the earliest time $C_{i}$ is reached, and define $h_{i}^{\pi}(s)=\mathbb{P}^{\pi}(\tau_{i}<\infty|s)$ to be the probability of reaching $C_{i}$ in finite time starting from state $s$ under $\pi$. We have then the following result.
@@
In general, for an irreducible stochastic transition matrix $P^{\pi}$ it is well-known (Seneta, 2006) that the only real-valued eigenvector associated to the eigenvalue $1$ is the unitary vector, and thus $r_{\lambda}$ is a solution for any $\lambda\in\mathbb{R}$.

### v3

@@ 6 Numerical Results
In this section, we present the numerical results of MR-NaS, and other algorithms, on various environments with different reward sets (results use $30$ seeds).
@@ Discussion.
The results for the first two settings are shown in Figures 3 and 4 (policy evaluation was performed using the MDP estimate $M_{t}$ at each time-step). MR-NaS achieves good accuracy on all environments. On the other hand, SF-NR and GVFExplorer have mixed performance. While SF-NR is not designed to optimize a behavior policy, we note that the exploration strategy used by GVFExplorer is similar to solving a problem akin to Equation 2, but neglects the forward equations when optimizing the behavior policy (see also Section C.3 for details). As a result, GVFExplorer tends to perform worse in environments where certain rewards are hard to obtain under a uniform policy. Lastly, we refer the reader to Section C.4 for more details, and to the README.md file in the supplementary material to reproduce the results.
@@ 7 Conclusions
In this work, we studied the problem of devising an exploration strategy for online multi-reward multi-policy evaluation, accommodating reward sets that are either finite or convex, potentially encompassing all possible rewards. Leveraging tools from Best Policy Identification, we derived an instance-dependent sample complexity lower bound for the $(\epsilon,\delta)$-PAC setting. Based on this bound, we extended MR-NaS (Russo & Vannella, 2024) to the policy-evaluation setting, and showed its asymptotic efficiency. Lastly, we compared MR-NaS against other adaptive exploration methods across various domains, demonstrating its efficiency.
@@ Appendix A Additional Notation
${\rm Alt}_{\pi,r}^{\epsilon}(M)$ ${\rm Alt}_{\pi,r}^{\epsilon}(M)=\{M_{r}^{\prime}:M\ll M_{r}^{\prime},\;\|V_{M_{r}}^{\pi}-V_{M_{r}^{\prime}}^{\pi}\|_{\infty}>2\epsilon\}$ Set of alternative models in $(\pi,r)$.
${\cal R}_{\pi}$ ${\cal R}_{\pi}\subseteq[0,1]^{S}$ Reward space considered for policy $\pi$.
${\cal R}_{\pi}^{\epsilon}$ ${\cal R}_{\pi}^{\epsilon}=\{r\in{\cal R}_{\pi}:{\rm Alt}_{\pi,r}^{\epsilon}(M)\neq\emptyset\}$ The subset of rewards of ${{\cal R}}_{\pi}$ for which the set of alternative models is not empty.
$\bf 1$} ${\bf 1}=\begin{bmatrix}1&\dots&1\end{bmatrix}$ all-ones vector.
$e_{i}$ $e_{i}=\begin{bmatrix}0&\dots&\underbrace{1}_{\hbox{i-th position}}&0&\cdots\end{bmatrix}$ $i$-th element of the canonical basis.
@@
$P_{s}^{\pi}$ $P_{s}^{\pi}=(P(s^{\prime}|s,\pi(s)))_{s^{\prime}\in{\cal S}}$ Vector of transition probabilities under $\pi$ in $s$.
@@
$G^{\pi}$ $G^{\pi}=(I-\gamma P^{\pi})^{-1},G^{\pi}\in\mathbb{R}^{S\times S}$ Discounted fundamental matrix.
[…]
@@
$\mathrm{sp}(V)$ $\max_{s}V(s)-\min_{s}V(s)$ Span of a value function.
$\Gamma^{\pi}(s)$ $\Gamma^{\pi}(s)=K^{\pi}(s)G^{\pi}$ Matrix used to express $\rho_{r}^{\pi}(s,s^{\prime})=e_{s^{\prime}}^{\top}\Gamma^{\pi}(s)r$.
$\mathrm{KL}_{P\|P^{\prime}}(s,a)$ $\mathrm{KL}\!\left(P(\cdot\mid s,a)\,\|\,P^{\prime}(\cdot\mid s,a)\right)$ KL between next-state kernels at $(s,a)$.
$\|q\|_{\rm TV}$ $\frac{1}{2}\|q\|_{1}$ Total variation norm for signed vectors $q$.
$d^{\pi}(s)$ $d^{\pi}=d^{\pi}P^{\pi},\;\sum_{s}d^{\pi}(s)=1$ Stationary state distribution under $\pi$.
$d^{\pi}(s,a)$ $d^{\pi}(s)\,\pi(a|s)$ Stationary state–action occupancy.
$\omega(s,a)$ $\omega\in\Delta({\cal S}\times{\cal A})$ Stationary state–action allocation (design variable).
$N_{t}(s,a)$ $\sum_{n=1}^{t}\mathbf{1}_{\{(s_{n},a_{n})=(s,a)\}}$ Visit count to $(s,a)$ up to time $t$.
@@ B.1 Alternative Models and Value Deviation
We begin by proving Proposition 4.1 and an additional necessary (minimax) condition for a model $M^{\prime}$ to be an alternative model.
B.1.1 Proof of Proposition 4.1 and an Additional Minimax Necessary Condition
@@
Sufficient condition. Let $s_{0},s_{1}\in{\cal S}$, and fix a reward $r\in{\cal R}$. In the following we omit the subscript $r$ for simplicity. Define a transition function $P^{{}^{\prime}}$ so that
@@
Letting $M=\sum_{i}(\gamma P^{{}^{\prime}\pi})^{i}e_{s_{0}}$, observe that $M_{s_{0}}=e_{s_{0}}^{\top}M$ represents the (unnormalized) stationary discounted probability of reaching $s_{0}$ starting from $s_{0}$
@@
Let $p_{s_{0}}^{\pi}=P(s_{0}|s_{0},\pi(s_{0}))$ . Then, we seek a value of $\delta\in(0,1)$ so that
@@
Necessary condition. As before, let $V_{M}^{\pi}$ be the value of $\pi$ in $M$ with a reward $r$ (we omit the subscript $r$ for simplicity), and similarly define $V_{M^{\prime}}^{\pi}$. If $M^{\prime}$ is confusing for some reward $r\in{\cal R}$, then we have that there exists a state $s_{0}$ such that $2\epsilon<|\Delta V^{\pi}(s_{0})|$, with $\Delta V^{\pi}(s)=V_{M}^{\pi}(s)-V_{M^{\prime}}^{\pi}(s)$. Note that the following inequalities hold for any $s$
$|\Delta V^{\pi}(s)|$ $\leq\gamma\left|P(s,\pi(s))^{\top}V_{M}^{\pi}-P^{\prime}(s,\pi(s))^{\top}V_{M^{\prime}}^{\pi}\right|,$
@@
In the following proof we provide an alternative minimax necessary condition for the existence of an alternative model for $(\pi,r)$. To derive the result, we use the following lemma.
@@
[…]
@@
Consider a reward $r$ and a target policy $\pi$. If $M^{\prime}$ is a confusing model, i.e. $V^{\pi}_{M^{\prime}}-V^{\pi}_{M}||_{\infty}>2\epsilon$, then there $\exists s$ such that $\mathbb{E}_{a\sim\pi(\cdot|s)}[{\rm KL}_{P|P^{\prime}}(s,a)]>\frac{2(1-\gamma)^{4}}{\gamma^{2}}\epsilon^{2}.$
@@
The proof relies on constructing an alternative MDP and a policy $\bar{\pi}$ such that we can transform the problem of evaluating two different transition functions into a problem of comparing two different policies.
@@
where $d_{\bar{M}_{\bar{r}}}^{\bar{\pi}}(s)$ denotes the visitation distribution induced by $\bar{\pi}$ in $\bar{M}_{\bar{r}}$ starting from $s$.
@@
$\frac{(1-\gamma)^{4}}{\gamma^{2}}\epsilon^{2}\leq\|\bar{\pi}(s)-\bar{\pi}^{\prime}(s)\|_{TV}^{2}\leq\frac{1}{2}{\rm KL}(\bar{\pi}(s),\bar{\pi}^{\prime}(s))\leq\frac{1}{2}\mathbb{E}_{a\sim\pi(\cdot|s)}\left[{\rm KL}_{P|P^{\prime}}(s,a)\right].$
@@ B.1.2 Additional Results on the Value Deviation $\rho$
${\rm sp}(V_{r}^{\pi})\coloneqq\max_{s^{\prime}}V_{r}^{\pi}(s^{\prime})-\min_{s}V_{r}^{\pi}(s),$ (11)
@@
For any reward vector $r\in[0,1]^{S}$, states $s,s^{\prime}\in{\cal S}$, we have: (I) $|\rho_{r}^{\pi}(s,s)|\leq 1$; (II) $|\rho_{r}^{\pi}(s,s^{\prime})|\leq{\rm sp}(V_{r}^{\pi})$ and (III) $|\rho_{r}^{\pi}(s,s^{\prime})-\rho_{r}^{\pi}(s^{\prime},s)|\leq|\Delta_{r}^{\pi}(s,s^{\prime})|+\Gamma_{s,s^{\prime}}^{\pi}{\rm sp}(V_{r}^{\pi}),$ where $\Delta_{r}^{\pi}(s,s^{\prime})\coloneqq r(s,\pi(s))-r(s^{\prime},\pi(s^{\prime}))$ and $\Gamma_{s,s^{\prime}}^{\pi}\coloneqq\frac{1+\gamma}{2}\|P(s,\pi(s))-P(s^{\prime},\pi(s^{\prime}))\|_{1}$.
@@
Then, define ${\cal Z}(s,s^{\prime})\coloneqq\{z\in{\cal S}:P(z|s^{\prime},\pi(s^{\prime}))\geq P(z|s,\pi(s))\}$ and
@@
The second statement stems from the simple fact that $V_{r}^{\pi}(s^{\prime})\leq-\min_{s}V_{r}^{\pi}(s)$ for any $s^{\prime}$. Hence:
@@
[…]
@@
Second part. For the second part of the lemma let $P_{s}^{\pi}=P(s,\pi(s))$ be the vector of transition probabilities in $s$ under action $\pi(s)$. We first prove that $V_{r}^{\pi}(s)\geq\lambda_{r}^{\pi}V_{r}^{\pi}(s^{\prime})\Rightarrow(P_{s}^{\pi})^{\top}V_{r}^{\pi}\geq\lambda_{r}^{\pi}(P_{s^{\prime}}^{\pi})^{\top}V_{r}^{\pi}$ . The proof is simple, and follows from the following inequalities that hold for any pair $(s,\hat{s})\in{\cal S}^{2}$:
@@
    We now prove a lower bound on $\rho_{r}^{\pi}(s,s^{\prime})$. The idea is to seek a value of $\eta\geq 0$ such that $\gamma W_{s^{\prime}}-W_{s}\geq(\gamma-1)W_{s}\eta$, which implies $\gamma W_{s^{\prime}}-W_{s}\geq-\eta$ (using that $W_{s}\geq-1/(1-\gamma)$). Then, we find $W_{s^{\prime}}\geq\frac{(\eta\gamma-\eta+1)}{\gamma}W_{s}$. Since $W_{s^{\prime}}\geq\lambda_{r}^{\pi}W_{s}$, we can set
@@
In the following, we prove Proposition 4.2, which characterizes the set of rewards for which $\max_{s}\|\rho_{r}^{\pi}(s)\|_{\infty}=0$. In the proof we define the following deviation matrix
@@
where $P^{\pi}$ is a $S\times S$ matrix satisfying $(P^{\pi})_{s,s^{\prime}}=P(s^{\prime}|s,\pi(s))$ (and $P_{s}^{\pi}=(P(s^{\prime}|s,\pi(s)))_{s^{\prime}\in{\cal S}}$). We consider the condition $\rho_{r}^{\pi}=0$ being identically 0, that is
@@ Conclusion.
The set of all vectors $r$ for which $\rho_{r}^{\pi}=0$ is precisely the set $\{\alpha{\bf 1}:\alpha\in\mathbb{R}\}$. Thus, $r$ must be a scalar multiple of $\bf 1$}. ∎
@@
Let ${\rm diag}(\rho_{r}^{\pi})=\begin{bmatrix}\rho_{r}^{\pi}(s_{1},s_{1})&\dots&\rho_{r}^{\pi}(s_{S},s_{S})\end{bmatrix}^{\top}$. We have that ${\rm diag}(\rho_{r}^{\pi})=(I-P^{\pi})(I-\gamma P^{\pi})^{-1}r.$ (13)
@@
We now consider a result that permits us to understand when $\max_{s}\rho_{r}^{\pi}(s,s)=0$ for any state $s$, which is the diagonal of $\rho_{r}^{\pi}$. In the result, we consider a partition of the set of recurrent states of $P^{\pi}$ into $m$ disjoint closed irreducible sets $C_{1},\dots,C_{m}$. Then, define $\tau_{i}=\inf\{t:s_{t}\in C_{i}\}$ be the hitting time of $C_{i}$, that is, the earliest time $C_{i}$ is reached, and define $h_{i}^{\pi}(s)=\mathbb{P}^{\pi}(\tau_{i}<\infty|s)$ to be the probability of reaching $C_{i}$ in finite time starting from state $s$ under $\pi$. We have then the following result.
@@
In general, for an irreducible stochastic transition matrix $P^{\pi}$ it is well-known (Seneta, 2006) that the only real-valued eigenvector associated to the eigenvalue $1$ is the all-ones vector, and thus $r_{\lambda}=\lambda\mathbf{1}$ is a solution for any $\lambda\in\mathbb{R}$.